### v1

@@ 4.2 Machine Learning Methods
To evaluate the predictive accuracy of the machining learning and linear regression models, we perform a validation process involving 1,000 random data splits. For each split, 80% of the observations are randomly selected as the training set for tuning and training the models, while the remaining 20% forms the test set for assessing predictive performance. The out-of-sample mean squared errors (MSE) are calculated for each model and split on the test set. The resulting average test MSEs across the 1,000 splits for the XGBoost, Random Forest, as well as the linear regression models selected by the step-wise procedure and LASSO, are presented in Table 4.2. We can see that the two linear regression models produce similar forecasting results, and the two machine learning models generate more accurate forecasts than their linear counterparts. XGBoost produces the lowest average test MSE across all models. Therefore, we select XGBoost as the preferred machine learning model for the subsequent analysis.
@@
4.3 Interpreting the Trained XGBoost Model
We now proceed to interpret the results of the XGBoost model fitted with the hyperparameter values presented in Table B.2 to the full sample. To achieve this, we employ feature importance and accumulated local effects (ALE) plots, both widely used tools for interpreting machine-learning-based models.
4.3.1 Feature Importance Plots
Figure 4.1 presents the relative importance of each feature in the trained XGBoost model, measured by its contribution to improving the goodness of fit of the model. The plot shows that expected loss is the most important variable for explaining CAT bond spreads, in line with the findings in existing studies and the linear regression models. In contrast to the linear regression models, it is interesting to note that PFL is also important in improving the predictive power of the XGBoost model even after including EL. This suggests that despite the high correlation between EL and PFL, the latter still contributes additional useful information in predicting CAT bond spreads. The linear regression model likely excludes PFL due to its nonlinear contribution to the spreads, a characteristic that XGBoost can capture.
Figure 4.1: Feature importance plot for the XGBoost model.
Besides EL and PFL, it is also worth considering the roles of other variables in the model. The Lane Financial LLC Synthetic RoL Index, $\mathrm{ROLX}$, ranks as the second most important variable in the XGBoost model. Besides, the Guy Carpenter indices, $\mathrm{GC.EU}$ and $\mathrm{GC.GLOB}$, designed to track the changes in reinsurance pricing for property catastrophe coverage in the European and global markets respectively, rank as the fifth and sixth most important. These observations are consistent with the fact that CAT bond investors expect a return that is commensurate with the risk they are taking, and this return is often benchmarked against the cost of traditional reinsurance. If reinsurance rates rise, investors demand higher yields from CAT bonds too. The pricing of reinsurance also reflects the market perception of risk, particularly for catastrophic events. Changes in reinsurance prices may signal a change in the perceived risk of catastrophic events, which affects CAT bond pricing as investors and issuers reassess the risk-return profile. The importance of the three indices—ROLX, GC.EU, and GC.GLOB—is also observed in the linear regression model.
$\mathrm{SIZE}$ ranks as the fourth most important variable in the XGBoost model, while this variable was excluded from the linear regression model selected with the step-wise procedure (c.f. Table 4.1). The exclusion of $\mathrm{SIZE}$ from the linear regression model could arise from the nonlinear relationship between $\mathrm{SIZE}$ and the bond spread and potential collinearity with other variables, which complicates the isolation of its unique contribution in a linear framework. XGBoost model can capture nonlinear relationships and interactions and is more robust to collinearity since it selects features based on their contribution to improving predictions across multiple decision trees. As a result, the XGBoost model recognizes the importance of $\mathrm{SIZE}$ even in the presence of correlated predictors. Finally, variables considered important in the linear regression model, such as $\mathrm{IG}$ and $\mathrm{GC.US}$, contribute minimally to the goodness of fit in the XGBoost model.
4.3.2 Accumulated Local Effect Plots
The feature importance plot provides an initial overview of the relative contribution of each predictor to the model. However, this plot alone does not reveal how each predictor affects the CAT bond spreads. To address this, we further explore the impact of the six most important predictors on CAT bond spreads by examining their accumulated local effect (ALE) plots, as shown in Figure 4.2. All ALE plots are centered to obtain an average of zero for enhanced interpretability. ALE plots are similar to the widely used partial dependence plots (PDP) but address the bias that arises in PDP when predictors are highly correlated (Apley and Zhu,, 2020). For instance, as EL and PFL are highly correlated, it is unlikely that a CAT bond would have a very high EL and very low PFL simultaneously, or vice versa. PDPs are not able to cope with such situations as they rely on marginal distributions of the predictors. Consequently, when evaluating the impact of EL on bond spreads, PDPs treat EL independently with all other predictors, and would consider unrealistic cases, for example, in which EL is high and PFL is very low. ALE plots rely on conditional distributions, and thus account for the correlation among predictors, leading to more accurate interpretations of the estimated model. Hence, the ALE plot is a more appropriate choice for our analysis considering the potential high correlations among predictors.
From Figure 4.2, nonlinearity is observed in the ALE plots. Expected loss shows the strongest impact on bond spreads. When EL is close to zero, the bond spread is approximately 350 bps lower than the average bond spread. As EL increases to 15%, the bond spread rises to approximately 1,250 bps higher than the average bond spread. For EL values below 5%, there is an almost linear relationship, where bond spreads increase with the expected loss. Beyond the 5% threshold, the relationship becomes less straightforward. A sharp increase in bond spread is observed as EL increases from 5% to 6.25%. Nevertheless, as EL progresses from 6.25% to 10%, the increase in bond spread is minimal. For EL values beyond 10%, the bond spread experiences a modest increase with EL. Caution should be taken in interpreting the section with large EL values exceeding 10%, due to the limited number of observations in this range.
The ALE plot for PFL indicates that higher probabilities of first loss are associated with increased bond spreads. This observation aligns with existing literature (Beer and Braun,, 2022) that identifies a positive correlation between PFL and bond spread. However, our interpretation differs significantly. In Beer and Braun, (2022), where only PFL is used as a predictor, the effect of EL is indirectly captured by PFL due to their high correlation. Consequently, the impact attributed to PFL in their model is a combined effect of both PFL and EL. In contrast, our XGBoost model incorporates both PFL and EL as predictors, with EL being identified as the most significant factor. Given their strong correlation, much of the information represented by PFL is also encompassed within EL. As a result, PFL contributes to the XGBoost model with information that is not already accounted for by EL. Therefore, the XGBoost model effectively utilizes all the information contained in EL and PFL. The impact of this additional information in PFL is captured in the ALE plot.
The impact of PFL on the bond spread is most significant at lower levels of PFL and diminishes to almost zero as PFL exceeds 5%. This indicates that at lower PFL values, investors are sensitive to changes in PFL and demand a higher spread for an increased PFL. However, as PFL reaches higher levels, further increases in PFL appear to only have a moderate impact on bond spread. A likely explanation for this observation is that when PFL is high, EL tends to be high as well due to their correlation. In such cases, investors may become less sensitive to small increases in PFL, shifting their focus more towards EL as the more influential factor in their decision-making process.
The impacts of ROLX and GC.EU on bond spreads also exhibit clear non-linear patterns. The impact of GC.GLOB on the bond spread is similar to but less significant than that of ROLX and GC.EU. Both ROLX and GC.EU have a modest influence on bond spread when their values are below 120 which signifies a relatively soft reinsurance market. In a soft market, there is generally an abundance of capital and a competitive environment. Investors might also perceive lower risk in this environment, resulting in a lower impact of changes in reinsurance market conditions on CAT bond spreads. In addition, CAT bonds can be viewed as alternatives to traditional reinsurance products. Reinsurance is still the dominant risk transfer mechanism for catastrophe risks and the risk transfer choice reflects the trade-off between the benefits and costs of reinsurance relative to CAT bond (Subramanian and Wang,, 2018). In a soft reinsurance market, where reinsurance capacity is ample and premiums are low, entities with hedging needs are more inclined to opt for reinsurance products rather than issuing CAT bonds. Small increases in reinsurance premiums are unlikely to divert these entities towards the CAT bond market, as the increased premiums remain affordable. Therefore, CAT bond spreads remain low and insensitive to reinsurance market conditions when the market is soft.
As the reinsurance market hardens, reflected by increased ROLX and the GC indices, there is tighter capital, higher demand for coverage, and increased perception of risk. Investors seek higher returns for assuming the same risks. The impact of ROLX and GC.EU on bond spread becomes more pronounced which is indicated by a steep increase in their corresponding ALE plots. Due to the limited capacity and high premiums in the reinsurance market, more entities with hedging demands turn to CAT bonds to transfer their risks. If reinsurance premiums continue to rise, more entities may shift towards the CAT bond market and thus drive up CAT bond spreads. This explains why CAT bond spreads are more sensitive to the reinsurance market conditions when the market is hard.
Our analysis also indicates that CAT bond spreads tend to increase with the issue size, particularly for issues smaller than 500 million US dollars. This observation is in contrast to the findings of Braun, (2016), who reported a negative impact of issue size on spreads. This difference results from the incorporation of GC indices in our model, which exhibit a moderate negative correlation (Pearson correlation of $-0.27$ for GC.EU and $-0.21$ for GC.GLOB) with the issue size. The negative correlation suggests that in a hardening market, characterized by tighter capital and higher perceived risks, CAT bond issue sizes tend to decrease. Therefore, issue size may serve as a weak indicator of market cycle conditions. In the absence of Guy Carpenter indices as predictors, as in the analysis by Braun, (2016), the observed negative impact of issue size on spreads could primarily reflect the effect of a hardening market. By incorporating GC indices as predictors in our model, we can distinguish the specific impact of issue size on bond spreads from the effects captured by the GC indices. The ALE plot suggests that in a hard market, where the issue size is often small, an increase in issue size is associated with an increase in bond spread.
Overall, the ALE plots highlight the presence of significant nonlinear relationships between key risk factors and CAT bond spreads. Failing to account for these nonlinear patterns could potentially lead to inaccurate estimations of bond spreads.
4.3.3 Second-Order Accumulated Local Effects Plots
Machine learning models like Random Forest and XGBoost naturally account for interaction effects between predictors, an advantage over linear regression where interactions must be explicitly defined. To visualize these interactions within the estimated XGBoost model, we utilize second-order ALE plots. While first-order ALE plots illustrate the “main” effect of individual predictors, second-order ALE plots reveal the additional joint impact of predictor pairs on the response variable.
As an illustration, Figure 4.3 displays second-order ALE plots for four predictor pairs: EL & ROLX, EL & SIZE, EL & GC.EU, and EL & GC.GLOB. The contour values on these plots indicate the strength of the joint impact. All four plots demonstrate clear interaction effects between the respective predictor pairs since the contour values vary significantly in each plot. The magnitude of these values is smaller compared to the ALE values depicted in Figure 4.2, suggesting that while these second-order effects are significant, they are generally smaller than the main effects of the individual predictors.
(a) EL (x-axis) and ROLX (y-axis)
(b) EL (x-axis) and SIZE (y-axis)
(c) EL (x-axis) and GC.EU (y-axis)
(d) EL (x-axis) and GC.GLOB (y-axis)
Figure 4.3: Second-order ALE plots of four selected predictor pairs.
We first examine Figure 3(a), which displays the interaction between EL and ROLX. At an EL value of 2%, an increase in ROLX from 120 to 140 does not change the spread. In contrast, at an EL of 8%, the same increase in ROLX results in an 80 bps decrease in the spread, thereby suggesting a negative interaction between EL and ROLX. It is important to note that the second-order ALE effect is exclusive of the main effect. To fully understand the impact of ROLX on the spread given the EL value, one must also consider the main effect of an increase in ROLX which is positive for ROLX within the range of 120–140. Specifically, an increase in ROLX from 120 to 140 results in approximately a 140 bps increase in the spread of a bond with a 2% EL, as illustrated in Figure 4.2. However, this increase translates to only a 60 bps (140 bps -- 80 bps) spread increase when EL is 8%. Therefore, for bonds with a higher EL, the impact of ROLX on the spread is still positive but weakened after combining the main and second-order effects.
Analysis of Figure 3(b) indicates that at an EL value of 2%, the increase in SIZE from 150 to 250 has a minimal effect on the spread. Nevertheless, the same increase in SIZE results in approximately 45 bps decrease in the spread when EL is 10%, suggesting a negative interaction between EL and SIZE. We have observed a positive main effect of SIZE on bond spread in Figure 4.2. The negative interaction between EL and SIZE suggests that an increase in issue size leads to a less pronounced rise, or even decrease, in the spread for bonds with a higher EL. Considering that the main effect of SIZE on the spread is relatively modest, as depicted in Figure 4.2, it is conceivable that the combined main and interaction effects of certain increases in SIZE could result in a negative impact on the spread for bonds with a high EL.
Figures 3(c) and 3(d) depict the interactions between EL and the Guy Carpenter indices, GC.EU and GC.GLOB, respectively. These interactions exhibit a complex pattern, neither consistently positive nor negative. For instance, at an EL value of 2%, increasing GC.EU from 140 to 160 does not affect the spread. At an EL of 4%, this same increase in GC.EU is associated with an increase in the spread. Interestingly, at an EL of 8%, an identical change in GC.EU results in a decrease in spread. The XGBoost model, with its tree-based structure, can capture these complex interactions, and thus offer more flexibility and higher prediction accuracy over models that assume linear relationships.
5 Probabilistic Analysis
In this section, we examine the probabilistic predictions produced by the linear regression model with normally distributed error terms and XGBoost employing the Jackknife+ prediction intervals introduced in Section 2.4.3. Section 5.1 presents the mean forecasts and prediction intervals generated by these two approaches. Section 5.2 further studies the conditional impact of expected loss and its interacting effects with a number of other factors.
5.1 Prediction Intervals
We first investigate the prediction intervals produced by the linear regression and the Jackknife+ Conformal Prediction with the XGBoost model. To this end, we randomly divide the full sample into a training set consisting of 80% of the data, and a test set with the remaining 20%. Both models are trained and tuned on the training set. Figure 5.1 presents the point forecasts (solid lines) and the corresponding 95% prediction intervals (shaded areas) produced by both models on the test data. The observed bond spreads (grey circles) are also depicted for reference and arranged in the order of their EL values. We observe that both models produce prediction intervals with a coverage close to the 95% target, while the coverage produced by XGBoost is slightly higher. Nevertheless, the average length of the prediction interval produced by XGBoost is approximately 13% smaller than that generated by the linear regression model. This suggests that the combination of Conformal Prediction with XGBoost yields more informative and precise prediction intervals.
(a) Linear regression
(b) XGBoost + Conformal Prediction
The results in Figure 5.1 are based on one random split, and thus could be subject to randomness in data splitting. To evaluate the robustness of this result, we calculate the average coverage and length of the 95% prediction intervals generated by the linear regression and XGBoost with Jackknife+ based on the 1,000 random splits of the full sample that are used in Table 4.2. The results are presented in Table 5.1. We can see that, similar to the results in Figure 5.1, the XGBoost with Jackknife+ model generates prediction intervals with average coverage closer to the desired 95% level and average length 8% smaller than the linear regression models.
@@ 5.2 Conditional Predictions
We further explore the conditional impact of several key risk factors using both linear regression and XGBoost models. Conditional impact refers to the effect of a single predictor on the predicted outcome, while keeping all other variables constant. To compute these conditional impacts, we initially calculate conditional predictions by varying the value of the risk factor under examination while holding all other factors at their median values for continuous variables or at the most common category for categorical variables. This process enables us to isolate and assess the impact of the specific risk factor. Furthermore, prediction intervals can be computed for these pseudo-observations, offering insights into the variation in conditional predictions.
Figure 5.2 depicts the conditional predictions of bond spreads across various EL values generated by both the linear regression model (red line) and XGBoost (blue line), and their corresponding 95% prediction intervals (shaded areas). Firstly, while the linear regression model predicts a linear impact of EL on CAT bond spreads, the XGBoost model exhibits nonlinear effects. Specifically, while increasing EL will in general lead to larger spreads, this effect is more pronounced for smaller EL values, and levels off when it exceeds 7%. Secondly, the linear regression model consistently yields higher spreads compared to XGBoost for all EL values. This observation suggests that the linear regression model tends to overestimate bond spreads, particularly for larger EL values. Importantly, the mean forecast produced by the linear regression model lies outside the 95% prediction interval generated by Conformal Prediction for the majority of EL values considered. This observation indicates that the overestimation by the linear model could be substantial.
@@
Figure 5.2 examines the conditional impact of EL on bond spread while all other risk factors are set at the median value or the most common category. We now further investigate the impact of EL under different reinsurance market conditions which are represented by ROLX values. Figure 5.3 displays the conditional impact of EL in three reinsurance market conditions. The blue line represents a hard reinsurance market, where the ROLX value is set as the average of the top 10% values, which is 144. The yellow line depicts a soft reinsurance market, where ROLX averages the bottom 10% value, 88. Finally, the red line corresponds to a normal market, with ROLX set at the median value, 116. Figures 3(a) and 3(b) present the predictions generated by the XGBoost and linear regression models, respectively, along with 95% prediction intervals for each conditional prediction.
Figure 3(a) demonstrates that using the XGBoost model, the shape of conditional prediction curves remains largely consistent across varying market conditions for different EL values, indicating a persistent nonlinear relationship between EL and spread. Nevertheless, the shifts of the conditional predictions across market conditions are not parallel. First, shifts from soft to normal market conditions result in only modest changes in conditional predictions for EL values below 7%, as evidenced by the close alignment of the red and yellow lines over this interval. This observation suggests that transitions from soft to normal market conditions have minor effects on the spreads of CAT bonds with lower EL values. However, a transition from normal to hard market conditions leads to a significant increase in conditional predictions for the same EL values. This indicates a lower investor sensitivity to the reinsurance market cycle in softer conditions, whereas in hard markets, characterized by constrained capacity and rising premiums, investors seek higher returns to offset the risks associated with CAT bonds, even those with lower EL values. Second, the upward shifts of bond spreads have comparable sizes for higher EL values when the reinsurance market transitions from soft to normal, and from normal to hard. This suggests that CAT bonds with higher EL values are sensitive to the tightening of reinsurance market conditions, even when the market is soft. These non-parallel changes of conditional impact curves across market conditions suggest an interaction between reinsurance market conditions and expected loss. In contrast, in Figure 3(b) where the linear regression model is applied, the conditional prediction curve shifts in parallel with changes in market conditions. The equal magnitude of shift, when market condition changes, reflects that the linear regression model only captures the linear effects of ROLX.
@@
Figure 5.4 illustrates the conditional impact of EL for bonds with different issue sizes. The blue line represents a high issuance volume, where the SIZE value is set as the average of the top 10% values, amounting to \$374.5 million. The yellow line depicts a low issuance volume, averaging the bottom 10% value, which is \$14.96 million. The red line corresponds to a median SIZE value of \$100 million. Two observations consistent with Figures 4.2 and 3(b) can be drawn. First, for small EL values, bonds with low issuance volumes generally have lower spreads compared to those with high or medium issuance volumes. This indicates that SIZE has a positive impact on bond spreads. However, this impact appears to be relatively minor. Second, when EL is large, SIZE exhibits a negative impact on spread. This is attributed to the negative interaction between SIZE and EL, the impact of which is stronger for high EL values, as discussed in Section 4.3.3.
@@ 6 Conclusion
In this study, we employ probabilistic machine learning techniques to analyze CAT bond pricing. Specifically, we use the XGBoost ensemble learning method to generate point forecasts of bond spreads, accompanied by Conformal Prediction to construct prediction intervals. This machine learning method has demonstrated improved prediction accuracy and more reliable prediction intervals compared to widely used linear regression models. It also provides insights into the complex relationships between various risk factors and CAT bond spreads and interaction effects among risk factors.
Our analysis utilizes a comprehensive dataset of primary market CAT bond transactions from January 1999 to March 2021. We begin with extensive linear regression models, and then proceed to feature selection through Lasso Regression and a stepwise approach. This process identifies expected loss (EL) as a primary predictor of CAT bond spreads, while another important risk factor, the probability of first loss (PFL), is excluded due to its strong correlation with EL. When employing the XGBoost model, interestingly, PFL is identified as the third most important factor. This indicates that, even with its correlation to EL, PFL offers valuable predictive information not captured by EL alone. Nonlinear impact on bond spread has been observed from factors such as EL, PFL, issue size, and market-wide indices representing reinsurance market conditions. Additionally, we uncover nontrivial interacting effects among these factors. Furthermore, we find that the XGBoost model, combined with the Conformal Prediction method, consistently produces not only more accurate point forecasts but also prediction intervals that are closer to the desired coverage level and exhibit smaller lengths.
This study advances our understanding of CAT bond pricing dynamics through machine learning techniques. These advancements offer tangible benefits to stakeholders in the CAT bond market, including investors, issuers, and policymakers. The improved understanding of risk factors and market sensitivities we provide is crucial for effective risk management and strategic decision-making. Future research could explore the nonlinear relationships revealed in this study, incorporate additional risk factors, and expand our dataset to further improve prediction accuracy. In addition, experimenting with alternative machine learning algorithms may offer new insights into the evolving landscape of CAT bond pricing. Finally, investigating the pricing dynamics of CAT bonds in the secondary market presents an intriguing avenue for future exploration.
Acknowledgement Hong Li is supported in part by funding from the Social Sciences and Humanities Research Council [430-2022-00401].

### v2

@@ 6.1 Predictive Accuracy and Intervals
The resulting average test MSEs across the 1,000 splits for the XGBoost and the linear regression models selected by the step-wise procedure and LASSO are presented in Table 6.1. For comparison purposes, we also include the test MSE for Random Forest. The two linear regression models produce similar forecasting results, while the two machine learning models demonstrate greater accuracy, generating lower MSEs than their linear counterparts. XGBoost produces the lowest average test MSE across all models.
@@
Since the linear regression models selected using the step-wise method and Lasso regularization lead to similar predictions, we only consider the model selected by step-wise method in the following analysis. The average coverage and length of the 95% prediction intervals generated by the linear regression and XGBoost with Jackknife+ based on the 1,000 random splits are presented in Table 6.2. The XGBoost model with Jackknife+ produces prediction intervals with an average coverage closer to the desired 95% level and an average length 8% smaller than those generated by the linear regression models. This suggests that the combination of Conformal Prediction with XGBoost yields more informative and precise prediction intervals.
@@
Figure 6.1 presents the point forecasts (solid lines) and the corresponding 95% prediction intervals (shaded areas) generated by both models for the test data based on a single split. The observed bond spreads (grey circles) are also depicted for reference and are arranged in order of their EL values. We observe that both models produce prediction intervals with a coverage close to the 95% target, with XGBoost achieving slightly higher coverage. Additionally, the average length of the prediction interval produced by XGBoost is approximately 13% smaller than that generated by the linear regression model. This reaffirms our earlier finding that integrating Conformal Prediction with XGBoost leads to more precise and informative prediction intervals.
(a) Linear regression
(b) XGBoost + Conformal Prediction
@@ 6.2 Conditional Predictions
We further exmaine the conditional impact of several key risk factors using both linear regression and XGBoost models. Conditional impact refers to the effect of a single predictor on the predicted outcome, while keeping all other variables constant. To compute these conditional impacts, we initially calculate conditional predictions by varying the value of the risk factor under examination while holding all other factors at their median values for continuous variables or at the most common category for categorical variables. This process allows us to isolate and assess the impact of the specific factor. Additionally, prediction intervals can be computed for these pseudo-observations, offering insights into the variation in conditional predictions.
Figure 6.2 depicts the conditional predictions of bond spreads across various EL values generated by both the linear regression model (red line) and XGBoost (blue line), along with their corresponding 95% prediction intervals (shaded areas). First, while the linear regression model predicts a linear impact of EL on CAT bond spreads, the XGBoost model exhibits nonlinear effects. Specifically, while increasing EL generally leads to larger spreads, this effect is more pronounced for smaller EL values and levels off when it exceeds 7%. Second, the linear regression model consistently yields higher spreads compared to XGBoost for all EL values. This observation suggests that the linear regression model tends to overestimate bond spreads, particularly for larger EL values. Notably, the mean forecast produced by the linear regression model lies outside the 95% prediction interval generated by Conformal Prediction for the majority of EL values considered. This finding indicates that the overestimation by the linear model could be substantial.
@@
Figure 6.2 examines the conditional impact of EL on bond spread while all other risk factors are set at the median values or the most common categories. We now further investigate the impact of EL under different reinsurance market conditions, represented by ROLX values. Figure 6.3 displays the conditional impact of EL across three reinsurance market conditions. The blue line represents a hard reinsurance market, where the ROLX value is set as the average of the top 10% values, which is 144. The yellow line depicts a soft reinsurance market, where ROLX averages the bottom 10% value, 88. Finally, the red line corresponds to a normal market, with ROLX set at the median value of 116. Figures 3(a) and 3(b) present the predictions generated by the XGBoost and linear regression models, respectively, along with 95% prediction intervals for each conditional prediction.
Figure 3(a) demonstrates that with the XGBoost model, the shape of conditional prediction curves remains largely consistent across varying market conditions for different EL values, indicating a persistent nonlinear relationship between EL and spreads. However, the shifts in the conditional predictions across market conditions are not parallel. First, shifts from soft to normal market conditions result in only modest changes in conditional predictions for EL values below 7%, as evidenced by the close alignment of the red and yellow lines over this interval. This observation suggests that transitions from soft to normal market conditions have minor effects on the spreads of CAT bonds with lower EL values. In contrast, a transition from normal to hard market conditions leads to a significant increase in conditional predictions for the same EL values. This indicates lower investor sensitivity to the reinsurance market cycle in softer conditions, whereas in hard markets, characterized by constrained capacity and rising premiums, investors seek higher returns to offset the risks associated with CAT bonds, even those with lower EL values. Second, the upward shifts of bond spreads are comparable in size for higher EL values when the reinsurance market transitions from soft to normal, and from normal to hard. This suggests that CAT bonds with higher EL values are sensitive to the tightening of reinsurance market conditions, even when the market is soft. These non-parallel changes in conditional impact curves across market conditions suggest an interaction between reinsurance market conditions and expected loss. In contrast, in Figure 3(b), where the linear regression model is applied, the conditional prediction curve shifts in parallel with changes in market conditions. The equal magnitude of shift when market conditions change reflects that the linear regression model only captures the linear effects of ROLX.
@@
Figure 6.4 illustrates the conditional impact of EL for bonds with different issue sizes. The blue line represents a high issuance volume, where the SIZE value is set at the average of the top 10% values, amounting to \$374.5 million. The yellow line depicts a low issuance volume, averaging the bottom 10% value, which is \$14.96 million. The red line corresponds to a median SIZE value of \$100 million. Two observations, consistent with Figures 5.2 and 3(b), can be made. First, for small EL values, bonds with low issuance volumes generally have lower spreads compared to those with high or medium issuance volumes, indicating that SIZE has a positive impact on bond spreads. However, this impact appears to be relatively minor. Second, when EL is large, SIZE exhibits a negative impact on spreads. This negative impact is attributed to the interaction between SIZE and EL, which is stronger for high EL values, as discussed in Section 5.3.
@@ 7 Conclusion
In this study, we employ probabilistic machine learning techniques to analyze CAT bond pricing. Specifically, we use the XGBoost ensemble learning method to generate point estimates of bond spreads, accompanied by Conformal Prediction to construct prediction intervals. This machine learning method provides insights into the complex relationships between various risk factors and CAT bond spreads, as well as the interaction effects among these factors. It has also demonstrated improved accuracy and more reliable prediction intervals compared to widely used linear regression models.
Our analysis utilizes a comprehensive dataset of primary market CAT bond transactions from January 1999 to March 2021. We begin with extensive linear regression models, followed by feature selection through Lasso Regression and a stepwise approach. This process identifies expected loss (EL) as a primary explanatory variable for CAT bond spreads, while another important factor, the probability of first loss (PFL), is excluded due to its strong correlation with EL. Interestingly, when employing the XGBoost model, PFL emerges as the third most important factor, indicating that despite its correlation with EL, PFL provides valuable information about CAT bond spreads not captured by EL alone. Nonlinear impacts on bond spreads have been observed from factors such as EL, PFL, issue size, and market-wide indices representing reinsurance market conditions. Furthermore, we uncover significant interaction effects among these factors. Finally, we find that the XGBoost model, combined with the Conformal Prediction method, consistently produces not only more accurate point forecasts but also prediction intervals that are closer to the desired coverage level and exhibit smaller lengths.
This study advances our understanding of CAT bond pricing dynamics through machine learning techniques. These advancements offer tangible benefits to stakeholders in the CAT bond market, including investors, issuers, and policymakers. The improved understanding of risk factors and market sensitivities we provide is crucial for effective risk management and strategic decision-making. Future research could experiment with alternative machine learning algorithms, which may offer new insights into the evolving landscape of CAT bond pricing. Additionally, investigating the pricing dynamics of CAT bonds in the secondary market presents an intriguing avenue for future exploration.